OpenStudy (anonymous):

if a number is subtracted from 3, the principal square root of this difference is equal to the number decreased by 1. find the number(s)?

OpenStudy (missinglink):

Let the unknown number be x. If we subtract 3 from x we have (x-3); we are told that the principal square root of this difference is equal to x - 1. Therefore sqrt(x-3) = x -1. We raise both sides of the equation to power 2 and we get x-3 = x^2 -2x + 1 x^2 -3x -2 = 0 (x-2)(x-1) = 0 So x is either 2 or 1.

OpenStudy (missinglink):

Nvm, it should be x^2 -3x + 4 = 0. So x = (3+7i)/2 or x = (3-7i)/2

OpenStudy (anonymous):

it was 2, :)

OpenStudy (anonymous):

thx